For two states defined by dV = 0 and dT  0, prove that the thermodynamic variables are connected through relation
(dv/dp)T(dp/dT)V(dT/dv)P= -1
1
Expert's answer
2020-04-15T10:26:22-0400

We know for the simple compressible substance,

"du=C_vdT+(\dfrac{du}{dv})_T dv"


"ds=\dfrac{C_v}{T}dT+\dfrac{1}{T}[(\dfrac{du}{dv})_T +P]dv"

Now, taking entropy as a function of temperature and volume,

"\dfrac{\partial s}{\partial T}_v=\dfrac{C_v}{T}"

Now, using maxwell relation,

"(\dfrac{\partial s}{\partial V})_T=(\dfrac{\partial P}{\partial T})_s=\dfrac{1}{T}[(\dfrac{\partial u}{\partial v})_T+P]"

From the above, we will get,

"(\dfrac{\partial u}{\partial v})_T=T(\dfrac{\partial P}{\partial T})_v-P"

Now, we know that PV=nRT

"(\dfrac{\partial P}{\partial T})_v=\dfrac{R}{V}"

"(\dfrac{\partial u}{\partial v})_T=T(\dfrac{R}{V})_v-P=P-P=0"

So,

"ds=\dfrac{C_p}{T}dT+[(\dfrac{dV}{dT})_P]dP"

From the above,

"(\dfrac{dP}{dT})_v=\dfrac{C_p-C_v}{T(\partial V/\partial T)_P}"


now, "\beta=\dfrac{1}{V}(\dfrac{\partial V}{\partial T})_P"


"k=-\dfrac{1}{V}(\dfrac{\partial V}{\partial T})_T"

Hence, from the above, we can write it as

"(\dfrac{\partial P}{\partial T})_v(\dfrac{\partial T}{\partial V})_P(\dfrac{\partial V}{\partial P})=-1"